OpenStudy (anonymous):

Can anyone please help to let me know if I am solving this the right way? f(x) = x^4 + 5, g(x) = x - 4 and h(x) = sqrt(x) f(g(h(x))) = h(x) = sqrt(x) = g(sqrt(x)) = sqrt(x-4)...

OpenStudy (anonymous):

I have to find f(g(h(x)))

OpenStudy (anonymous):

no your plugging in gx into hx but ur suppose to plug in hx into gx

OpenStudy (anonymous):

yeah there is a lot of variables and so I am getting confused

OpenStudy (anonymous):

yeah its reasonable but u had the right idea just plug in the function of h(x) for every x you see in g(x)

OpenStudy (anonymous):

h(x)= sqrt(x) = g(x) = x - 4?

OpenStudy (anonymous):

its asking you to plug in hx into gx or (g(hx)) so... sqrtx-4 now try the next part

OpenStudy (anonymous):

(sqrtx)-4 the 4 is outside the sqrt

OpenStudy (anonymous):

so you keep working backwards starting with h(x) and then g(h(x)) and then f(g(h(x)))?

OpenStudy (campbell_st):

if \[h(x) = \sqrt{x}\] then \[g(h(x)) = g(\sqrt{x}) = \sqrt{x} - 4 \] so for f(g(h(x))) in f(x) replace x with \[\sqrt{x} - 4\] hope it helps

OpenStudy (anonymous):

where does the x^4+5 come in for f(x)?

OpenStudy (campbell_st):

ok... so replacing x with \[\sqrt{x} - 4\] its \[f((g(h(x))) = f(\sqrt{x} -4) = (\sqrt{x} - 4)^4 + 5\]

OpenStudy (anonymous):

there is a lot of substation in a problem like this

OpenStudy (anonymous):

I did some problems before this that were just f(g(x)) but it got more complicated with f(g(h(x)))

OpenStudy (campbell_st):

its a tough problem.... probably at the top end of questions...

OpenStudy (anonymous):

ok, so I put in f(g(h(x))) = (sqrt(x-4)^4 + 5) and still got it incorrect

OpenStudy (anonymous):

is that because we need to simplify?

OpenStudy (campbell_st):

I'd say so.... you need to expand it.... I'd suggest you look at wolfram alpha and get it so simplify it

OpenStudy (anonymous):

I will put it into wolfram alpha

OpenStudy (campbell_st):

yep just go simplify (sqrt(x) - 4)^4 + 5 then enter it should work

OpenStudy (anonymous):

ok wolfram alpha came back with (x-4)^2 + 5

OpenStudy (anonymous):

still not right... looking at the problem again

OpenStudy (campbell_st):

well this is what I have \[x^2 - 16\sqrt{x^3} + 96x -256\sqrt{x} + 261\]

OpenStudy (anonymous):

ok let me plug that in

OpenStudy (anonymous):

that was it... can you walk me through it now? that would be extremely helpful

OpenStudy (anonymous):

I have an exam on tuesday

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well you have \[(\sqrt{x} - 4)^2 \times (\sqrt{x} -4)^2 + 5\] so I distributed and got \[(x - 8\sqrt{x} + 16) \times (x - 8\sqrt{x} + 16)\] then its a very tedious process of \[x(x - 8\sqrt{x} + 16) - 8\sqrt{x}(x - 8\sqrt{x} + 16)+ 16(x - 8\sqrt{x} + 16) + 5\] etc etc ... hope it makes sense...

OpenStudy (anonymous):

wow... that is a lot of work

OpenStudy (anonymous):

is there not a faster way to do the problem? I am bound to make and error doing all that work

OpenStudy (campbell_st):

well you could use binomial expansion.... the coefficients are 1 4 6 4 1 from pascals triangle \[(\sqrt{x} - 4) = 1 \times (x^{\frac{1}{2}})^4 \times (-4)^0 + 4 \times (x^{\frac{1}{2}})^3 \times (-4)^1 + ... \] the powers of x decrease and the powers of (-4) increase

OpenStudy (anonymous):

oh my instructor mentioned pascal's triangle in class but I don't even know that concept so that is out of the question...

OpenStudy (anonymous):

never been introduced to pascal's triangle

OpenStudy (anonymous):

maybe for now the long way is my only choice

OpenStudy (campbell_st):

I'd say so... I'd expect its a very unfair examination question... just because of the amout of time needed to get a solution

OpenStudy (anonymous):

yeah, I'd agree with that... so for now in my homework I will put the tedious work down so at least it shows I made an effort to get to the solution